Suppose a firm has a retention ratio of 23 percent and net income of $4.3 million. How much does it pay out in dividends? ( Enter in dollars not in millions).

Dividend Amount _______________
Answer #1

First we will calculate the dividend payout ratio as below:

Dividend payout ratio = 1 - Retention ratio

Dividend payout ratio = 1 - 0.23 = 0.77 or 77%

Now,

Dividend amount = Net income * Dividend payout ratio

Dividend amount = $4300000 * 77% = $3311000